OpenStudy (anonymous):

how do you solve this? .5^n = .2

OpenStudy (anonymous):

i have no idea...

OpenStudy (anonymous):

@Luis_Rivera can you please explain it a bit more clearly?

OpenStudy (anonymous):

@Luis_Rivera

OpenStudy (anonymous):

\[0.5^{n}=0.2\]

OpenStudy (anonymous):

\[\ln 0.5^{n} = \ln 0.2\]

OpenStudy (anonymous):

\[n \ln 0.5 = \ln 0.2\]

OpenStudy (anonymous):

\[n= \frac{ Ln(0.5) }{Ln(0.2) }\]

OpenStudy (anonymous):

Thanks guys! I got it!